This is my question,
Find the equation of the line parallel to the line:
2x-y+4=0, that passes through the point (5,2). Write the equation in point slope and slopoe intercept forms..
So far I have
2x-y+4=0
y=-2+2x
Am I on the Right path ???
Thank you

Answer by Alan3354(69443)

Find the equation of the line parallel to the line:
2x-y+4=0, that passes through the point (5,2). Write the equation in point slope and slopoe intercept forms..
So far I have
2x-y+4=0
y=-2+2x
Am I on the Right path ???
------------
It's y = 2x + 4, but that doesn't affect the problem.
Parallel lines have the same slope.
---
Use y-y1 = m*(x-x1) where m = slope and (x1,y1) is the point.